# Convolution goes higher-order: a biologically inspired mechanism empowers image classification

### v1

@@ 3 Experiments & Results
We evaluated the performance of our model on multiple datasets, including a synthetic one consisting of textures with higher order correlations and four widely-used benchmarks: MNIST, FashionMNIST, CIFAR-10, CIFAR-100 and Imagenette. This section details our experimental setup and findings.
@@ 3.1 Synthetic Dataset: Structured Visual Textures
To rigorously test the sensitivity of our model to higher-order correlations, we generated a synthetic dataset of structured visual textures based on the work of Victor & Conte (2012) (see Figure 2). These textures allow precise control over the type and intensity of contained correlations.
To generate the textures, we utilized a custom software library (Piasini et al., 2021) implementing the method developed by Victor & Conte (2012). This process involves sampling from a distribution of binary textures with specific multipoint correlation probabilities while maximizing entropy. The correlation intensity is parameterized by parity counts of white or black pixels within 1-, 2-, 3-, or 4-pixel tiles (termed gliders, Figure 2)). Notably, two-point and three-point gliders can produce multiple distinct multipoint correlations based on their spatial configurations. Two-point correlations can arise from horizontal, vertical, or oblique gliders, while three-point correlations can give rise to L patterns with various orientations. Our synthetic dataset comprises 2000 training images, 1000 validation images, and 2000 testing images. All reported results are based on the testing set.
@@
Table 1: Accuracy and parameters of CNN and HoCNN models on Texture classification
@@ 3.2 Benchmark Datasets
3.2.1 MNIST, FashionMNIST, CIFAR-10, CIFAR-100
To further validate our approach and ensure its generalizability, we conducted extensive experiments on four well-established benchmark datasets: MNIST, FashionMNIST, CIFAR-10, and CIFAR-100.
We focused on these relatively small datasets to allow for comprehensive testing and analysis. To ensure the robustness of our results and account for the inherent variability in neural network training, we conducted multiple independent runs for each experiment. Specifically, we initialized and trained each model with 50 different random seeds. This approach provides a more reliable estimate of model performance by mitigating the effects of random initializations.
@@ 3.2.1 MNIST, FashionMNIST, CIFAR-10, CIFAR-100
Training Details. For all, experiments previously introduced experiments, we used AdamW optimizer (Loshchilov, 2017) with learning rate 0.001, weight decay 5e-4, batch size 64, and cross-entropy loss. Input images were normalized using z-score standardization, without any additional data augmentation.
3.2.2 Imagenette
To investigate the scalability and real-world applicability of our approach, we conducted experiments with deeper architectures and a more challenging dataset. Specifically, we implemented Higher-order Convolution in a ResNet-18 (He et al., 2016) architecture (HoResNet-18) and evaluated it on Imagenette (Howard, 2020), a subsampled dataset from ImageNet containing 10 classes. The HoResNet-18 achieved better test accuracy compared to the standard ResNet-18 (see Table 2), demonstrating a performance improvement of over 1 percentage point on this challenging dataset.
The HoResNet-18 follows a similar structure to the standard ResNet-18, with a key modification in the first stage, where higher-order residual blocks are used. The remaining stages utilize standard residual blocks. This hybrid approach allows for the incorporation of higher-order convolutions while maintaining the overall structure of the ResNet architecture. The parameter count for both models is provided in Table 2, while architectural details are presented in the Appendix A.3.1. Additionally, our HoResNet-18 architecture achieves superior performance compared to a ResNet-18 adapted VOneNet (Dapello et al., 2020), a biologically-inspired model, while utilizing fewer parameters, as detailed in Appendix A.4.
These results provide strong evidence for the efficacy of our higher-order convolution approach, demonstrating successful scaling to deeper architectures and more complex datasets. By consistently outperforming the baseline across diverse benchmarks and multiple random initializations, our model exhibits robust improvements in image classification performance beyond simple tasks.
Training Details. For Imagenette specifically, images were normalized using mean = [0.5, 0.5, 0.5] and standard deviation = [0.5, 0.5, 0.5], in order to align with VOneNet (Dapello et al., 2020) preprocessing. All the other training details remain consistent with Subsection 3.2.1.
Figure 3: (A) Standard CNN architecture for Image Classification Benchmarks, the Conv Block represents the concatenation of Conv + Batch Norm. + ReLU + Pooling layers (B) Higher-Order CNN architecture: the first layer is split in three branches, representing the three orders of the Volterra expansion. After computing them separately, they’re added together (mimicking once again the Volterra series expansion. Similarly, the HoConv Block represents the concatenation of HoConv + Batch Norm. + ReLU + Pooling layers. For both networks, the number of kernels depend on the classification task.
Table 2: Accuracy comparison (with $\pm$ $\sigma$, standard deviation) of CNN and HoCNN models across image classification benchmarks, averaged over 50 seeds.
Dataset Accuracy (%) $\pm$ $\sigma$ Params
@@ 4 Image Statistics Sensitivity and Neural Representations
To directly quantify HoCNN’s sensitivity to higher-order image statistics, we conducted a systematic perturbation analysis on the CIFAR-10 test dataset using both architectures (50 pretrained models each), while keeping the network trained on the unperturbed images. We generated synthetic textures with varying orders of statistical correlations (1-, 2-, 3-, and 4-point) and interpolated them with test images at different intensity (I) levels (imgperturbed = (1-I)$\times$×imgCIFAR-10 + I×$\times$×imgperturbation , with I $\in[0.05,0.20]$). While HoCNN achieves superior performance on CIFAR-10 test set images (see Table 2), it shows increased vulnerability to higher-order statistical perturbations. The performance degradation becomes progressively more pronounced as we move from lower to higher-order correlations, with the gap widening at higher perturbation intensities (see Figure 4 A and B for I = 0.12). For instance, with third-order perturbations at I = 0.12, HoCNN’s accuracy drops by 78.0% compared to CNN’s 73.7% (complete results in Appendix A.6).
This enhanced sensitivity to statistical perturbations suggests HoCNN’s increased capacity to leverage higher-order patterns. To better understand how this statistical sensitivity translates into improved classification performance, we examined the model’s internal representations using Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) on CIFAR-10, analyzing activations from 100 test images (10 per class) averaged across 50 model realizations. The Conv and HoConv blocks exhibit distinct representational geometries (Figure 4A & B), with the latter showing more pronounced class-specific patterns. Analysis of pairwise distance distributions (Figure 4D) reveals that HoConv representations are significantly more dispersed in the high-dimensional space, indicating its ability to capture more diverse and discriminative features. Detailed analysis of individual order components and layer-wise correlations is presented in Appendix A.7.
@@
Figure 4: Perturbation Analysis and Neural Representations. (A) An exemplar test image (CIFAR-10) and perturbed examples with different 1, 2, 3, 4-point correlations statistics and common intensity (I = 0.12). (B) Normalized (wrt unperturbed CIFAR-10) accuracy for Convolutional (Blue) vs Higher-order (Orange) networks: performances of Higher-order network are systematically worse for more structured perturbations at fixed intensity (I = 0.12). (C) Representational Dissimilarity Matrix (RDM) for baseline Convolutional (Conv) Block (Conv layer, Batch Norm, ReLU, Pooling). (D) RDM for the Higher-order Convolutional (HoConv) block. (E) Log Ratio between the two RDMs, capturing different representational geometries between the two blocks. (F) An alternative view of the HoConv block effectiveness versus the Conv block, comparing the pairwise distance distribution across entries of each RDM. The orange distribution, representing the HoConv block, is significantly shifted to the right, supporting the enhanced image classification performances observed in Figure 5 and Table 2.
5 Discussion
@@
Our study introduces a novel approach to image classification by extending convolutional neural networks with higher-order learnable convolutions, inspired by complex nonlinear biological visual processing. Our findings indicate that expansion beyond the 4th order is unnecessary, a conclusion supported by both theoretical insights and empirical evidence. Koenderink & van Doorn (2018) analysis of natural image statistics reveals that the quadric, cubic and quartic power dominates approximately 63%, 35%, and 2% of the pixels, respectively. Our experimental results align with this distribution, showing no significant performance gains beyond the 3rd order in benchmark tasks and the 4th in structured visual textures. Through systematic perturbation analysis, we validate this alignment by isolating the contributions of specific image statistics to model performance. While adversarial perturbations are typically employed to assess and improve network robustness (Szegedy, 2013; Goodfellow et al., 2014) or analyze network behavior (Carter et al., 2019; Santurkar et al., 2019), we leverage them to probe our model’s sensitivity to distinct orders of image statistics, demonstrating how different orders of convolution process specific aspects of the input.
@@
Our approach to structured visual textures relates our findings to current mechanistic models in computational neuroscience, particularly those focusing on the Drosophila visual system, which is known to be sensitive and selective to different 3-point correlations (Fitzgerald & Clark, 2015), (Clark et al., 2014). This connection strengthens the relevance of our model and highlights its potential to provide insight into visual processing in both biological and artificial systems.
The consistent outperformance of our Higher-order CNN (HoCNN) across various datasets demonstrates the effectiveness of incorporating these biologically inspired computations. Representational Similarity Analysis (RSA) reveals distinct geometries, providing evidence that our model processes visual information differently from standard CNNs, suggesting different strategies for encoding and processing visual information.
@@
Future research directions include exploring deeper architectures with multiple higher-order convolutional layers, which presents challenges in balancing increasing nonlinearity across layers. To a certain extent, the scaling technique we employed (see Section 2) revealed better results compared to the case when it was not used. While computational complexity represents a current limitation, we envision future improvements through low-rank tensor decomposition methods. These techniques, which have proven successful in transformer architectures (Hu et al., 2021), could maintain the benefits of higher-order interactions while significantly reducing computational overhead.
@@
In conclusion, our higher-order CNN represents a promising direction for more effective and biologically inspired computer vision models. By exploiting visual patterns and relationships more efficiently, even in shallow architectures, our approach opens new possibilities for advancing the field. Future research should focus on understanding the interplay between different orders of nonlinearity, especially for spatiotemporal inputs (i.e. videos), on exploring more complex architectures, and investigating potential applications beyond image classification, such as object detection or semantic segmentation.

### v2

@@ 3 Experiments & Results
We evaluated our model on a synthetic texture dataset with higher-order correlations and multiple standard benchmarks: MNIST, FashionMNIST, CIFAR-10, CIFAR-100, and ImageNet. Additional results on Imagenette and fine-grained classification (CUB-200-2011) are provided in Appendix A.6 and A.7.
@@ 3.1 Synthetic Dataset: Structured Visual Textures
To test our model’s sensitivity to higher-order correlations, we generated a synthetic dataset of structured visual textures based on the work of Victor and Conte [37] (see Figure 2). These textures allow precise control over the type and intensity of contained correlations. To generate the textures, we utilized a custom software library [38] implementing the method developed by Victor and Conte [37]. This process involves sampling from a distribution of binary textures with specific multipoint correlation probabilities while maximizing entropy. The correlation intensity is parameterized by parity counts of white or black pixels within 1-, 2-, 3-, or 4-pixel tiles (termed gliders, Figure 2)). Notably, two-point and three-point gliders create distinct correlations through different spatial configurations: horizontal, vertical, or oblique for two-point, and L patterns with various orientations for three-point correlations. Our synthetic dataset comprises 2000 training images, 1000 validation images, and 2000 testing images. All reported results are based on the test set.
@@
Table 1: CNN and HoCNN performances on Texture classification
@@ 3.2 Benchmark Datasets
MNIST, FashionMNIST, CIFAR-10, CIFAR-100
To further validate our approach and ensure its generalizability, we conducted extensive experiments on four well-established benchmark datasets: MNIST, FashionMNIST, CIFAR-10, and CIFAR-100. We focused on these relatively small datasets to allow for comprehensive testing and analysis. To ensure the robustness of our results and account for the inherent variability in neural network training, we conducted multiple independent runs for each experiment. Specifically, we initialized and trained each model with 50 different random seeds. This approach provides a more reliable estimate of model performance by mitigating the effects of random initializations.
@@
Training Details. For all experiments, we used AdamW optimizer [39] with learning rate 0.001, weight decay 5e-4, batch size 64, cross-entropy loss and early stopping. Images were normalized using z-score standardization without data augmentation. Experiments used an NVIDIA RTX 4080 Ti GPU, with training times of 10-15 minutes for MNIST/FashionMNIST and 20-30 minutes for CIFAR-10/CIFAR-100 (HoCNN requiring  1.5-2× longer). Complexity analysis is in Appendix A.2.
IMAGENET
To investigate the scalability and real-world applicability of our approach, we implemented Higher-order Convolution in a ResNet-18 [4] architecture (HoResNet-18) and evaluated it on ImageNet [40]. To ensure statistical robustness, we conducted 5 independent training runs with different random seeds. HoResNet-18 achieved 69.47% $\pm$ 0.09% test accuracy compared to ResNet-18’s 68.62% $\pm$ 0.08%, demonstrating a statistically significant improvement of 0.85% (see Table 2). Results on Imagenette and fine-grained classification (CUB-200-2011) are reported in Appendix A.6 and A.7.
HoResNet-18 maintains the standard ResNet-18 structure, but uses higher-order residual blocks in its first stage and standard blocks thereafter. This hybrid design preserves ResNet’s architecture while incorporating higher-order convolutions. Parameter counts are shown in Table 2, with architectural details in Appendix A.9.1. These results provide strong evidence for the efficacy of our higher-order convolution approach, demonstrating successful scaling to deeper architectures and large-scale datasets. To further investigate scalability, we conducted extensive experiments with deeper ResNet architectures (ResNet-34 and ResNet-50), confirming that performance improvements persist across model depths (see detailed results in Appendix A.4).
Training Details. For ImageNet, we used standard practices: normalization with mean = [0.485, 0.456, 0.406] and standard deviation = [0.229, 0.224, 0.225]; Stochastic Gradient Descent with momentum = 0.9 as optimizer. We conducted 5 independent runs with different random seeds to ensure statistical robustness. The ImageNet experiments were run on an NVIDIA A100 GPU, requiring approximately 2-3 days per run. The detailed analysis of computational requirements and FLOP measurements for different orders of convolution can be found in Appendix A.2.
Figure 3: (A) Comparison of standard convolution (Conv) and (B) higher-order convolution (HoConv) blocks.
Table 2: Accuracy (with $\sigma$, standard deviation over multiple initializations) of CNN and HoCNN models across image classification benchmarks. MNIST, FashionMNIST, CIFAR-10, CIFAR-100: 50 runs; ImageNet: 5 runs.
Dataset Accuracy (%) $\pm$ $\sigma$ # Params
@@ 4 Image Statistics Sensitivity and Neural Representations
[…]
@@
Figure 4: Perturbation Analysis and Neural Representations. (A) An exemplar test image (CIFAR-10) and perturbed examples with different 1, 2, 3, 4-point correlations statistics and common intensity (I = 0.12). (B) Normalized (wrt unperturbed CIFAR-10) accuracy for Convolutional (Blue) vs Higher-order (Orange) networks: performances of Higher-order network are systematically worse for more structured perturbations at fixed intensity (I = 0.12). (C) Representational Dissimilarity Matrix (RDM) for baseline Convolutional Block (Conv layer, Batch Norm, ReLU, Pooling). (D) RDM for the Higher-order Convolutional block. (E) Log Ratio between the two RDMs, capturing different representational geometries between the two blocks.
5 Discussion and Limitations
@@
Our study introduces a novel approach to image classification by extending convolutional neural networks with higher-order learnable convolutions, inspired by complex nonlinear biological visual processing. Our findings indicate that expansion beyond the 4th order is unnecessary, a conclusion supported by both theoretical insights and empirical evidence. Koenderink and van Doorn [1]’s analysis of natural image statistics reveals that the quadratic, cubic and quartic power dominates approximately 63%, 35%, and 2% of the pixels, respectively. Our experimental results align with this distribution, showing no significant performance gains beyond the 3rd order in benchmark tasks and the 4th in structured visual textures. Through systematic perturbation analysis, we validate this alignment by isolating the contributions of specific image statistics to model performance.
@@
Our approach to structured visual textures connects to mechanistic models in computational neuroscience, particularly those studying the Drosophila visual system’s sensitivity to 3-point correlations [19, 20]. This connection strengthens the relevance of our model and highlights its potential to provide insight into visual processing in both biological and artificial systems. The consistent outperformance of our HoCNN across various datasets demonstrates the effectiveness of incorporating these biologically inspired computations. RSA reveals distinct geometries, providing evidence that our model processes visual information differently from standard CNNs.
5.1 Limitations
Despite the demonstrated benefits, our higher-order convolution approach has several limitations worth acknowledging. The primary challenge is computational complexity, as detailed in Appendix A.2. Our FLOP analysis reveals that 2nd and 3rd order operations require approximately 5× and 18× more computation than standard convolutions, respectively. While we mitigate this by strategic placement of higher-order operations in early network layers, this computational overhead represents a significant trade-off for resource-constrained applications. The increasing parameter count with higher-order expansions also introduces scalability challenges. Although our implementation leverages kernel symmetry to reduce parameters (e.g., reducing the 3rd order parameter count by 4.42×), memory requirements still increase substantially with order. This limits practical applications beyond 4th order expansions, though our experiments indicate minimal benefits beyond this point.
Additionally, our approach shows increased vulnerability to higher-order statistical perturbations (Section 4), suggesting a trade-off between enhanced feature extraction capability and certain forms of robustness. We also observed that placement of higher-order layers significantly impacts performance (Appendix A.3), with deeper placement leading to performance degradation. Finally, our current implementation does not employ low-rank approximations or other efficiency techniques commonly used in modern architectures, representing an opportunity for future optimization.
5.2 Future Directions
An important consideration is how our higher-order convolution approach relates to Vision Transformers [42], which have emerged as powerful alternatives to CNNs. While both capture feature interactions, they operate on fundamentally different principles. Our higher-order convolutions explicitly model multiplicative interactions between neighboring pixels through polynomial expansions, whereas transformers employ bilinear operations with global attention. From a mathematical perspective, attention mechanisms compute pairwise interactions through the query-key product ($QK^{T}$), which are inherently bilinear, while our approach extends to trilinear and higher-order multiplicative terms within local receptive fields. These differences suggest that higher-order convolutions and vision transformers may be complementary rather than competing approaches, potentially offering different advantages for capturing local complex patterns versus global relationships.
A particularly promising direction involves hybrid architectures that leverage both mechanisms: higher-order convolutions could process local complex patterns and higher-order correlations in early layers, while attention mechanisms capture long-range dependencies and global context in deeper layers. This combination could exploit the strengths of both approaches—efficient local feature extraction through higher-order polynomial expansions and flexible global reasoning through self-attention. Investigating such architectures could reveal whether these mechanisms provide complementary or redundant information, and identify optimal integration strategies for different vision tasks.
@@ 5.2 Future Directions
Beyond architectural hybridization, future research directions include exploring deeper architectures with multiple higher-order convolutional layers and investigating efficiency improvements through low-rank tensor decomposition methods. These techniques, which have proven successful in transformer architectures [43], could maintain the benefits of higher-order interactions while significantly reducing computational overhead. Additionally, extending our approach to other computer vision tasks beyond image classification, such as object detection and dense prediction tasks, represents an important avenue for validating the generalizability of higher-order convolutions.
@@
In conclusion, our higher-order CNN represents a promising direction for more effective and biologically inspired computer vision models. By exploiting visual patterns and relationships more efficiently, even in shallow architectures, our approach opens new possibilities for advancing the field. Future research should focus on understanding the interplay between different orders of nonlinearity, their relationship with attention mechanisms, and investigating applications across diverse vision tasks.